### v1

@@ 4.3 Lazy Multiscale
Let $q$ denote the convergence rate of an algorithm with iterate convergence on a loss function $\tilde{\mathcal{L}}$. Under the setting of Section˜4.1, the lazy multiscale method satisfies
$\left\lVert x_{1}^{K_{1}}-x_{1}^{*}\right\rVert$ $\leq\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert(q)^{K_{S}}\prod_{s=1}^{S-1}(1+(q)^{K_{s}})$
$\qquad+\ \tfrac{1}{2}L_{f^{*}}\left\lvert u-\ell\right\rvert\sum_{s=1}^{S-1}\frac{1}{\sqrt{2^{S-s}}}(q)^{K_{s}}\prod_{j=1}^{s-1}\left(1+(q)^{K_{j}}\right).$
When using constant iterations $K_{s}=K$ at each scale, this reduces to
@@
$\qquad+\ \tfrac{1}{2}L_{f^{*}}\left\lvert u-\ell\right\rvert\frac{d(K)\left(\sqrt{2^{S}}\left(d(K)+1\right)^{S}-\sqrt{2}\left(d(K)+1\right)\right)}{\sqrt{2^{S}}\left(d(K)+1\right)\left(\sqrt{2}d(K)+\sqrt{2}-1\right)}$
where $d(K)=(q)^{K}\in(0,1)$.
@@
See LABEL:sec-lazy-multiscale-descent-error-proof.
@@ 5 Relation Between the Discretized and Continuous Problems
We assume $t\in\mathbb{R}^{I}$ forms a uniform grid on the interval $\ell,u]$ with spacing $\Delta t=t[{i+1}]-t[i]=\frac{u-\ell}{I-1}$.
Lemma 5.1 (Distance Between Piecewise Linear Functions).
@@
$\|\hat{f}-\hat{g}\|_{2}^{2}\leq\tfrac{2}{3}\Delta t\|x-y\|_{2}^{2}.$
@@
See LABEL:sec-piecewise-function-approximation-proof.
Lemma˜5.3 compares favourably with existing results. De Boor [de_boor_practical_2001, Ch. III, Eq. 17] establishes the bound
@@
De Boor showed that the error vanishes for continuous functions (not necessarily Lipschitz), though possibly a slower rate than Lemma˜5.3 [de_boor_practical_2001, Ch. III, Eq. 18]. Kunoth et al. [kunoth_splines_2018, Sec. 1.5, Th. 18] generalizes the approximation error to functions in Sobolev spaces, achieving for differentiable Lipschitz functions a rate comparable to Lemma˜5.3:
@@
for some constant $K>0$.
We cannot expect a better rate because Lipschitz functions are differentiable almost everywhere [heinonen_lectures_2004]. For Lipschitz functions (possibly not differentiable), the derivative, where it exists, satisfies $\left\lvert f^{\prime}(x)\right\rvert\leq L_{f}$, giving $\left\lVert f^{\prime}\right\rVert_{2}^{2}\leq(u-\ell)L_{f}^{2}$. Thus, Lemma˜5.3 explicitly calculates the constant $K=2/15$ in Eq.˜9.
Theorem 5.5 (Connection Between the Continuous and Discrete Problems).
Let $f:[\ell,u]\to\mathbb{R}$ be $L_{f}$-Lipschitz with discretization $x^{*}\in\mathbb{R}^{I}$. Let $x\in\mathbb{R}^{I}$, and construct the corresponding piecewise linear function $\hat{f}_{x}$ (Definition˜2.4). For $\epsilon>0$, if
$I>C\epsilon^{-1}+1\quad\text{and}\quad\|x-x^{*}\|_{2}^{2}<D\epsilon$
then
@@ 6 Comparison with Projected Gradient Descent
Lemma˜2.7 establishes iterate convergence (cf. Definition˜2.6) for projected gradient descent. We use this to derive expected convergence bounds for three approaches: projected gradient descent at scale $s=1$ (Theorem˜6.1), greedy multiscale descent (Theorem˜6.4), and lazy multiscale descent (Theorem˜6.14). After presenting these results, we demonstrate that both multiscale variants achieve tighter error bounds with lower computational cost when the problem size and Lipschitz constant are sufficiently large.
@@
If we iterate projected gradient descent $K$ times, where
$K\geq\frac{\log(1/\epsilon)+{\log(2^{S}+2)}/{2}}{-\log q},$
then the expected error is bounded as
$\mathbb{E}\left\lVert x_{1}^{k}-x_{1}^{*}\right\rVert_{2}\leq\epsilon.$
@@
Assume the same setting as Theorem˜6.1, but use the greedy multiscale descent method starting at scale $s=S$. At the coarsest scale, the expected initial error satisfies
$\mathbb{E}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}\leq 2.$
Performing $K_{s}$ iterations at each scale yields the expected final error
(11) $\mathbb{E}\,\|x_{1}^{K_{1}}-x_{1}^{*}\|_{2}\leq(q)^{K_{1}}\sqrt{2^{S+1}}\left((q)^{r_{S}}+\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{{2^{S+2}}}\sum_{s=1}^{S-1}2^{s}(q)^{r_{s}}\right)$}
@@
See LABEL:sec-expected-greedy-convergence-proof.
Determining the iteration count required to achieve a specified accuracy is less direct for multiscale methods than for projected gradient descent. Moreover, a fair comparison requires accounting for computational cost: an iteration at coarse scale $S$ requires fewer floating-point operations than an iteration at the finest scale $s=1$. We therefore configure each iteration of multiscale to simultaneously achieve lower total cost than fine-scale-only optimization while obtaining a tighter expected error bound. Specifically, we require Eq.˜11 to be smaller than Eq.˜10. This requires costing multiscale in terms of one projected gradient descent iteration at the finest scale $s=1$.
@@
Suppose projected gradient descent with $K$ iterations at the finest scale $s=1$ has total cost $C_{\mathrm{GD}}=AK$, where $A>0$ is the cost per iteration at that scale. Assume the cost of one projected gradient descent iteration scales polynomially with problem size as $C_{s}=C(I_{s})^{p}$ for $I_{s}=2^{S-s+1}+1$ points and constants $C>0$ and $p\geq 1$. Then greedy multiscale descent with $K_{s}$ iterations at scale $s$ has total cost $C_{\mathrm{GM}}$ that satisfies
$C_{\mathrm{GM}}\leq A\sum_{s=1}^{S}K_{s}\left(\frac{3}{5}\right)^{s-1}.$
@@
See LABEL:sec-cost-of-gd-vs-ms-proof.
The factor $3/5$ is the ratio of the number of points at the two coarsest scales: $I_{S}/I_{S-1}=3/5$.
The assumption that iteration cost scales as $CI^{p}$ for $C>0$ and $p\geq 1$ holds for common projects and gradient computations. For example, projection of a vector $y\in\mathbb{R}^{I}$ onto the sphere $\mathcal{S}^{I-1}$, given by
[…]
scales linearly in II ($p=1$). The least-squares gradient update for $A\in\mathbb{R}^{I\times I}$,
[…]
scales quadratically in II ($p=2$).
Lemma˜6.6 establishes a selection of iterations counts $K_{s}$ at each scale such that greedy multiscale is cheaper than projected gradient descent.
Corollary 6.8 (Conditions for Greedy Multiscale to be Cheaper).
Consider projected gradient descent at scale $s=1$ with $K$ iterations. Greedy multiscale with $K_{s}$ iterations at scale $s$ achieves lower cost under either iteration plan:
  • $K_{s}=\left\lceil 2K/5\right\rceil-1$ iterations at every scale; or
  • $K_{s}=1$ iterations at every scale except the finest scale, where $K_{1}=K-2$.
Proof 6.9.
This follows from Lemma˜6.6 combined with bounds on the sums
$\sum_{s=1}^{S}\left(\frac{3}{5}\right)^{s-1}=\frac{1-(3/5)^{S}}{1-(3/5)}<5/2$
and
$\sum_{s=2}^{S}\left(\frac{3}{5}\right)^{s-1}=\frac{3/5-(3/5)^{S}}{1-(3/5)}<3/2<2.$
The analysis in Corollary˜6.8 ignores the cost of interpolation and memory allocation when changing iterate dimensions between scales. These costs become negligible as the iteration counts increase. Empirical results in Section˜7 confirm that greedy multiscale remains cheaper in practice.
@@
Proof 6.11.
See LABEL:sec-greedy-cost-proof.
@@
Suppose projected gradient descent with $K$ iterations has total cost $C_{\mathrm{GD}}=AK$, where $A>0$ is the cost per iteration at the finest scale $s=1$. Under the polynomial cost scaling assumption $C_{s}=\Theta(I_{s}^{p})$ for $I_{s}=2^{S-s+1}+1$ points and $p\geq 1$, lazy multiscale descent with $K_{s}$ iterations at scale $s$ has total cost $C_{\mathrm{LM}}$ that satisfies
$C_{\mathrm{LM}}\leq A\left(\sum_{s=1}^{S-1}2^{-s}K_{s}+2^{-S}3K_{S}\right).$
With $\left\lceil\frac{4}{5}K\right\rceil-1$ iterations at each scale,
$C_{\mathrm{LM}}<C_{\mathrm{GD}}.$
@@
See LABEL:sec-cost-of-gd-vs-lazy-ms-proof
Theorem 6.14 (Expected Lazy Multiscale Convergence).
Assume the same setting as Theorem˜6.1, but use the lazy multiscale starting at coarsest scale $S$ with $I=2^{S}+1$ points at the finest scale. The expected initial error satisfies
$\mathbb{E}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}=2.$
Performing constant $K_{s}=K$ iterations at each scale yields the expected final error
$\mathbb{E}\left\lVert x_{1}^{K}-x_{1}^{*}\right\rVert_{2}$
$\leq d(K)\left(2(d(K)+1)^{S-1}+\frac{L_{f^{*}}}{2}\left\lvert u-\ell\right\rvert\frac{\left(\sqrt{2^{S}}\left(d(K)+1\right)^{S}-\sqrt{2}\left(d(K)+1\right)\right)}{\sqrt{2^{S}}\left(d(K)+1\right)\left(\sqrt{2}d(K)+\sqrt{2}-1\right)}\right),$
Proof 6.15.
See LABEL:sec-expected-lazy-convergence-proof.
@@
$S\geq\frac{\log\left(\frac{q^{-K/5-1}}{1+q^{K^{\prime}}}\left(2+\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{2(\sqrt{2}(1+q^{K^{\prime}})-1)}\right)\right)}{\log\left(\frac{\sqrt{2}}{1+q^{K^{\prime}}}\right)},$
@@
Proof 6.17.
See LABEL:sec-lazy-cost-proof.
Corollaries˜6.10 and 6.16 reveal several insights. Both multiscale variants require a minimum number of scales $S$ (equivalently a minimum sized problem $I$) for the multiscale optimization to outperform single scale methods. Theorem˜5.5 establishes that arbitrarily small error between continuous and discretized solutions requires arbitrarily large problem size $I$. Both corollaries show that the required $I$ increases with the Lipschitz constant $L_{f^{*}}$ and the domain size $\left\lvert u-\ell\right\rvert$, consistent with Theorem˜5.5. A subtle requirement is that the base algorithm cannot have iterative convergence with $q=0$, which would enable solving the discretized problem in one iteration, and thus eliminate the benefit of multiscale.
Corollaries˜6.10 and 6.16 differ in their constraints on the convergence rate $q$. Greedy multiscale with one iteration per scale requires convergence rate $q\in(0,1/2)$ because interpolation error must remain smaller than the algorithmic progress at each scale. Lazy multiscale, on the other hand, allows $q\in(0,1)$ but requires a minimum of $K^{\prime}>\log_{q}(\sqrt{2}-1)$ iterations at each scale to sufficiently control error accumulation. This is a mild requirement: two iterations per scale suffice for lazy multiscale to improve on the base algorithm when $0<q<(\sqrt{2}-1)^{1/2}\approx 0.6436$, and both algorithms require arbitrarily many iterations to drive the final error to zero.
@@ 7 Numerical Experiments and Benchmarks
The project-gradient multiscale and single-scale algorithms are implemented in the Julia package BlockTensorFactorization.jl [Richardson_BlockTensorFactorization_jl]: the function factorize implements projected gradient descent for the Tucker-1 decomposition problem at a single scale, and multiscale_factorize implements the greedy multiscale method (Algorithm˜2). All experiments ran on an Intel Core i7-1185G7 with 32GB of RAM, without parallelization, matching the hardware used for the example experiment in Section˜2.3.
@@ 7.1 Density Demixing as Optimization
Graham et. al. [graham_tracing_2025] provides a full treatment of this formulation which we summarize here. Additional details are given in LABEL:additional-tensor-details-and-reference.
Given $I$ mixtures $y_{i}:\mathcal{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}_{+}$ of $R$ source probability density functions $b_{r}:\mathcal{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}_{+}$,
$y_{i}(x)=\textstyle\sum_{r\in[R]}a_{i,r}b_{r}(x),$
we seek to recover the sources $\{b_{r}\}$ and their mixing coefficients $a_{i,r}$. We assume $R<I$ and that the source densities $\{b_{r}\}$ are linearly independent, which ensures this is a well-posed problem.

### v2

@@ 4.3 Lazy Multiscale
Assume the same setup as Theorem˜4.9. The lazy multiscale method in Algorithm˜1 returns iterate $x_{1}^{K_{1}}$ satisfying
$\left\lVert x_{1}^{K_{1}}-x_{1}^{*}\right\rVert\leq\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert d(K_{S})p(S)+\tfrac{1}{2}L_{f^{*}}\left\lvert u-\ell\right\rvert\sum\nolimits_{s=1}^{S-1}\frac{1}{\sqrt{2^{S-s}}}d(K_{s})p(s),$
where $d(K)=(q)^{K}\in(0,1)$ and $p(s)=\prod_{j=1}^{s-1}(1+(q)^{K_{j}})$. When using constant iterations $K_{s}=K$ at each scale, this reduces to
@@
$\qquad+\ \tfrac{1}{2}L_{f^{*}}\left\lvert u-\ell\right\rvert d(K)\frac{\sqrt{2^{S}}\left(d(K)+1\right)^{S}-\sqrt{2}\left(d(K)+1\right)}{\sqrt{2^{S}}\left(d(K)+1\right)\left(\sqrt{2}d(K)+\sqrt{2}-1\right)}.$
@@ 5 Relation Between the Discretized and Continuous Problems
Lemma 5.1 (Piecewise Linear Function Distance).
@@
$\|\hat{f}-\hat{g}\|_{2}^{2}\leq\tfrac{2}{3}\Delta t\|x-y\|_{2}^{2},$
with equality holding when the end points are equal: $x[1]=y[1]$ and $x[I]=y[I]$.
@@
Lemma˜5.3 compares favourably with existing results. De Boor [7, Ch. III, Eq. 17] establishes the bound
@@
De Boor showed the error vanishes for continuous functions (not necessarily Lipschitz), at a slower rate than Lemma˜5.3 [7, Ch. III, Eq. 18]. Kunoth et al. [15, Sec. 1.5, Th. 18] generalizes the approximation error to functions in Sobolev spaces, achieving for differentiable Lipschitz functions a rate comparable to Lemma˜5.3:
@@
for some constant $K>0$. We should not expect a better rate because Lipschitz functions are differentiable almost everywhere [13]. For Lipschitz functions, the derivative, where it exists, satisfies $\left\lvert f^{\prime}(x)\right\rvert\leq L_{f}$, giving $\left\lVert f^{\prime}\right\rVert_{2}^{2}\leq(u-\ell)L_{f}^{2}$. Thus, Lemma˜5.3 explicitly calculates the constant $K=2/15$ in Eq.˜8.
Theorem 5.5 (Continuous Problem Connection).
Let $f:[\ell,u]\to\mathbb{R}$ be $L_{f}$-Lipschitz with discretization $x^{*}\in\mathbb{R}^{I}$. Let $x\in\mathbb{R}^{I}$, and construct the corresponding piecewise linear function $\hat{f}_{x}$ (Definition˜2.4). For $\epsilon>0$, if $I>C\epsilon^{-1}+1$ and $\|x-x^{*}\|_{2}^{2}<D\epsilon$ then
@@ 6 Comparison with Projected Gradient Descent
Lemma˜2.7 establishes iterate convergence (Definition˜2.6) for projected gradient descent. Using this, we derive expected convergence bounds for three approaches: projected gradient descent at scale $s=1$ (Theorem˜6.1), greedy multiscale descent (Theorem˜6.4), and lazy multiscale descent (Theorem˜6.12). We demonstrate that both multiscale variants achieve tighter error bounds with lower computational cost when the problem size is sufficiently large.
@@
If we iterate projected gradient descent $K$ times, where $K\geq\left(\log(1/\epsilon)+{\log(2^{S}+2)}/{2}\right)/(-\log q),$ then the expected error is bounded as $\mathbb{E}\left\lVert x_{1}^{k}-x_{1}^{*}\right\rVert_{2}\leq\epsilon.$
@@
Assume the same setting as Theorem˜6.1, but use the greedy multiscale descent method starting at scale $s=S$. Performing $K_{s}$ iterations at each scale yields the expected final error
(10) $\mathbb{E}\,\|x_{1}^{K_{1}}-x_{1}^{*}\|_{2}\leq(q)^{K_{1}}\sqrt{2^{S+1}}\left((q)^{r_{S}}+\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{{2^{S+2}}}\sum\nolimits_{s=1}^{S-1}2^{s}(q)^{r_{s}}\right)$}
@@
The result follows from Theorems˜6.1 and 4.9 with the initial error bound $\mathbb{E}\left[\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert\right]\leq(q)^{0}\sqrt{2^{1}+2}=2$ starting at the coarsest scale with $I=3$ points.
Determining the iteration count required to achieve a specified accuracy is less direct for multiscale methods than for projected gradient descent. Moreover, a fair comparison requires accounting for computational cost: an iteration at coarse scale $S$ requires fewer floating-point operations than an iteration at the finest scale $s=1$. We therefore configure each iteration of multiscale to simultaneously achieve lower total cost than fine-scale-only optimization while obtaining a tighter expected error bound. Specifically, we require Eq.˜10 to be smaller than Eq.˜9. This requires costing multiscale in terms of one projected gradient descent iteration at the finest scale $s=1$.
@@
Suppose projected gradient descent with $K$ iterations at the finest scale $s=1$ has total cost $C_{\mathrm{GD}}=CK$, where $C>0$ is the cost per iteration at that scale. Assume the cost $C$ of one projected gradient descent iteration scales polynomially $C=\Theta(I^{p})$ with the problem size $I$ for some constant $p\geq 1$. Then greedy multiscale descent with $K_{s}$ iterations at scale $s$ has total cost $C_{\mathrm{GM}}$ that satisfies $C_{\mathrm{GM}}\leq C\sum_{s=1}^{S}K_{s}\left(\frac{3}{5}\right)^{s-1}.$
With $K_{s}=1$ at every scale except the finest, where $K_{1}=K-2$, $C_{\mathrm{GM}}<C_{\mathrm{GD}}$.
@@
The assumption that iteration cost scales as $\Theta(I^{p})$ for $p\geq 1$ holds for common projects and gradient computations. For example, the Euclidean projection of a vector $y\in\mathbb{R}^{I}$ onto the sphere $\mathcal{S}^{I-1}$ scales linearly in II ($p=1$). The least-squares gradient update $x\leftarrow x-A^{\top}(Ax-y)$ for $A\in\mathbb{R}^{I\times I}$ scales quadratically in II ($p=2$).
The analysis in Lemma˜6.6 ignores the cost of interpolation and memory allocation when changing iterate dimensions between scales. These costs become negligible as the iteration counts increase. Empirical results in Figs.˜3 and 7 confirm that greedy multiscale remains cheaper in practice.
@@
Proof 6.9.
@@
Suppose the same set up as Lemma˜6.6. Lazy multiscale descent with $K_{s}$ iterations at scale $s$ has a total cost $C_{\mathrm{LM}}$ satisfying
$C_{\mathrm{LM}}\leq C\left(\textstyle\sum_{s=1}^{S-1}2^{-s}K_{s}+2^{-S}3K_{S}\right).$
With $K_{s}=\left\lceil\frac{4}{5}K\right\rceil-1$ iterations at each scale, $C_{\mathrm{LM}}<C_{\mathrm{GD}}.$
Proof 6.11.
Theorem 6.12 (Expected Lazy Multiscale Convergence).
Assume the same setting as Theorem˜6.1, but use the lazy multiscale starting at coarsest scale $S$ with $I=2^{S}+1$ points at the finest scale. Performing constant $K_{s}=K$ iterations at each scale yields the expected final error
(11) $\mathbb{E}\left\lVert x_{1}^{K}-x_{1}^{*}\right\rVert_{2}$
$\leq d(K)\left(2(d(K)+1)^{S-1}+\frac{L_{f^{*}}}{2}\left\lvert u-\ell\right\rvert\frac{\sqrt{2^{S}}\left(d(K)+1\right)^{S}-\sqrt{2}\left(d(K)+1\right)}{\sqrt{2^{S}}\left(d(K)+1\right)\left(\sqrt{2}d(K)+\sqrt{2}-1\right)}\right),$
@@
The proof is similar to Theorem˜6.4 by using Theorem˜4.11.
@@
$S\geq{\log\left(\frac{q^{-K/5-1}}{1+q^{K^{\prime}}}\left(2+\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{2(\sqrt{2}(1+q^{K^{\prime}})-1)}\right)\right)}\Big/{\log\left(\frac{\sqrt{2}}{1+q^{K^{\prime}}}\right)},$
@@
Proof 6.15.
Corollaries˜6.8 and 6.14 reveal several insights. Both multiscale variants require a minimum number of scales $S$ (equivalently a minimum sized problem $I$) for the multiscale optimization to outperform single scale methods. Theorem˜5.5 establishes that arbitrarily small error between continuous and discretized solutions requires arbitrarily large problem size $I$. Both corollaries show that the required $I$ increases with the Lipschitz constant $L_{f^{*}}$ and the domain size $\left\lvert u-\ell\right\rvert$, consistent with Theorem˜5.5. A subtle requirement is that the base algorithm cannot have iterative convergence with $q=0$, which would enable solving the discretized problem in one iteration, and thus eliminate the benefit of multiscale.
Corollaries˜6.8 and 6.14 differ in their constraints on the convergence rate $q$. Greedy multiscale with one iteration per scale requires convergence rate $q\in(0,1/2)$ because interpolation error must remain smaller than the algorithmic progress at each scale. Lazy multiscale, on the other hand, allows $q\in(0,1)$ but requires a minimum of $K^{\prime}>\log_{q}(\sqrt{2}-1)$ iterations at each scale to sufficiently control error accumulation. This is a mild requirement: two iterations per scale suffice for lazy multiscale to improve on the base algorithm when $0<q<(\sqrt{2}-1)^{1/2}\approx 0.6436$, and both algorithms require arbitrarily many iterations to drive the final error to zero.
@@ 7 Numerical Experiments and Benchmarks
The projected-gradient multiscale and single-scale algorithms are implemented in the Julia package BlockTensorFactorization.jl [20]: the function factorize implements projected gradient descent for the Tucker-1 decomposition problem at a single scale, and multiscale_factorize implements the greedy multiscale method (Algorithm˜1). All experiments ran on an Intel Core i7-1185G7 with 32GB of RAM, without parallelization, matching the hardware used for the example experiment in Section˜2.3.
@@ 7.1 Density Demixing as Optimization
Graham et. al. [11] provides a full treatment of this formulation which we summarize here. Additional details are given in LABEL:additional-tensor-details-and-reference.
Given $I$ mixtures $y_{i}:\mathcal{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}_{+}$ of $R$ source probability density functions $b_{r}:\mathcal{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}_{+}$, $y_{i}(x)=\textstyle\sum_{r\in[R]}a_{i,r}b_{r}(x),$ we seek to recover the sources $\{b_{r}\}$ and their mixing coefficients $a_{i,r}$. We assume $R<I$ and that the source densities $\{b_{r}\}$ are linearly independent, which ensures this is a well-posed problem.